OpenStudy (anonymous):

x+4/x-1 divided by x^2+x/x-1=

OpenStudy (kc_kennylau):

remember \(\Large\frac ab\div\frac cd=\frac ab\times\frac dc\) :)

OpenStudy (anonymous):

I just can not rap my head around where to start??? I am in week 4 MAT/!!& and it is hard for me

OpenStudy (kc_kennylau):

\[\frac{x+4}{x-1}\div\frac{x^2+x}{x-1}\\=\frac{x+4}{x-1}\times\frac{x-1}{x^2+x}\]

OpenStudy (anonymous):

? x^3+4x-1/x^2 ?

OpenStudy (kc_kennylau):

\[\frac{x+4}{x-1}\div\frac{x^2+x}{x-1}\\=\frac{x+4}{x-1}\times\frac{x-1}{x^2+x}\\=\frac{(x+4)(x-1)}{(x-1)(x^2+x)}\]

OpenStudy (anonymous):

I think I got that Thank you so much I as not reversing the C/d and multiplying I was trying to divide the whole.

OpenStudy (kc_kennylau):

no problem :) tell me your answer?

OpenStudy (anonymous):

No my bad redo :) x+4/x(x+1) !! Got it Finally Thanks